Why is the exact value of sin(inverse-1) -1 = -pi/2?
Answer by stanbon(75887)

Why is the exact value of sin(inverse-1) -1 = -pi/2?
----
sin^-1(-1) = -pi/2
---
sin^-1(x) means "the angle whose sine is x".
------------
So your problem says "the angle whose sine is -1 is -pi/2",
and this is a true statement. That is, the sin of (-pi/2)
is -1.
================
Cheers,
Stan H.
=============